OpenStudy (ksaimouli):

3cosx+9=7 slove for x saifoooo help

OpenStudy (ksaimouli):

131.8 another one

OpenStudy (ksaimouli):

plx 73 plz

OpenStudy (anonymous):

\[\color{red}{x}\color{blue}{\text{saifoo}}=\cos^{-1}(-\frac{2}{3})\]

myininaya (myininaya):

I told you how to get the solutions \[x=\cos^{-1}(\frac{-2}{3})+2n \pi\] Please ask me what you don't understand about how I got this? by the way n is an integer

OpenStudy (anonymous):

looks like 131.8 degrees if you are working in degrees

OpenStudy (anonymous):

hello myininaya!

myininaya (myininaya):

if you want degrees do +360 n

myininaya (myininaya):

and put your calc in degree mode for approximations

myininaya (myininaya):

hey sat

OpenStudy (ksaimouli):

how to get 228.19

OpenStudy (anonymous):

i am making truth tables. nothing could be more boring, but i have to do it t t t f f t f f arrghhh

OpenStudy (ksaimouli):

is i have to add 180 hhhhhhhhhh i dont know plz help

OpenStudy (anonymous):

subtract 131.8 from 360

OpenStudy (anonymous):

let me draw a picture

OpenStudy (ksaimouli):

ok

OpenStudy (anonymous):



myininaya (myininaya):

cos(x)=cos(-x)

OpenStudy (anonymous):

both those places on the circle have the same first coordinate, that is the same cosine. the first one you can get from the calculator, it is131.8 the other one you get from subtracting 131.8 from 360

myininaya (myininaya):

\[\cos(131.8)=\cos(-131.8)=\cos(-131.8+360)\]

OpenStudy (anonymous):

@myininaya is that a five o-clock shadow you're sporting?

OpenStudy (ksaimouli):

ok thx so ican sub 360 for everythink like thses problems